The length of a rectangle is 12. What is the width if the perimeter is 36.
Found 2 solutions by stanbon, Cromlix:
Answer by stanbon(75887)

The length of a rectangle is 12. What is the width if the perimeter is 36.
-----
P = 2(L + W)
36 = 2(12 + W)
18 = 12 + W
Width = 6
==============
Cheers,
Stan H.
==============

Answer by Cromlix(4381)

Perimeter = 2*length + 2*width
36 = 2(12) + 2(width)
36 = 24 = 2*width
2 * width = 36 - 24
2* width = 12
width = 12/2
width = 6.
Hope this helps.
:-)